OpenStudy (babynini):

f(x)=x^3, a=8 Find (f^-1)'(a)

OpenStudy (babynini):

well, f'(x) = 3x^2 f(0)=0

OpenStudy (tkhunny):

You're supposed to know this: \(\dfrac{d}{dx}f^{-1}(x) = \dfrac{1}{f'\left(f^{-1}(x)\right)}\)

OpenStudy (babynini):

hmm Yes. I know this. I'm just not sure how to apply it xD I find f^-1(x) first?

OpenStudy (anonymous):

no not really

OpenStudy (anonymous):

i mean you can if you want, but you don't have to since \[\dfrac{d}{dx}f^{-1}(x) = \dfrac{1}{f'\left(f^{-1}(x)\right)}\] you know \[\dfrac{d}{dx}f^{-1}(8) = \dfrac{1}{f'\left(f^{-1}(8)\right)}\]

OpenStudy (anonymous):

you can find \(f^{-1}(8)\) easily by solving \[x^3=8\] fror \(x\)

OpenStudy (babynini):

x=2

OpenStudy (babynini):

So then we have \[\frac{ 1 }{ f'(2) }=\frac{ 1 }{ 12 }\]

zepdrix (zepdrix):

yay good job

OpenStudy (babynini):

Yaay. Thanks @satellite73 :)